OpenStudy (anonymous):

2 (cos^2 (2 t) + sin^2 (2 t)) I do not remember where to go next to simplify but I am trying for cos^2+sin^2=1

OpenStudy (anonymous):

or i mean to ask what role does 2t take on

OpenStudy (anonymous):

should i just use double angle?

OpenStudy (anonymous):

i mean power reducing...

OpenStudy (anonymous):

imagine 2t as x, it's the angle; so cos^2 (x) + sin^2(x) = 1, then 2(1) = 2 ^_^

OpenStudy (anonymous):

i did the power reducing and got same thing, but thanks for that tip!

OpenStudy (anonymous):

lol, np :)

OpenStudy (anonymous):

I believe you do have to use the double angle formula.

OpenStudy (anonymous):

Use double angle formula since you cant really use (sinx)^2+(cosx)^2=1 on this example.

OpenStudy (nowhereman):

That is nonsense! Of course you can use\[\sin^2 x + \cos^2 x = 1\] because you can plug in x = 2t without a problem.

OpenStudy (anonymous):

why do you need the double angle formula? that makes stuff so complicated, we already know that cos^2(x) + sin^2(x) = 1 x acts as an angle and it can be anything, 2t, w, z, c whateva form, it's still an angle

OpenStudy (anonymous):

true dat

OpenStudy (anonymous):

hmm... interesting

OpenStudy (anonymous):

write it down on paper and you'll see what I mean @rg ^_^

OpenStudy (anonymous):

yes, she is so right, I mean, every time I write it down on paper everything looks much more understandable, dont try to do everything in your head, unless ofcourse, you are doing maths just to increase your brains processing power... anyway.. it is a good idea to write it down...

OpenStudy (anonymous):

lol andy ^_^